# Risk-Averse Total-Reward Reinforcement Learning

### v1

@@ 1 Introduction
Risk-averse reinforcement learning (RL) is an essential framework for practical and high-stakes applications, such as self-driving, robotic surgery, healthcare, and finance greenberg2022efficient ; fei2021exponential ; urpi2021risk ; hau2024q ; suevar ; zhang2021mean ; hau2023entropic ; Kastner2023 ; Marthe2023a ; lam2022risk ; li2022quantile ; bauerle2022markov ; pan2019risk ; mazouchi2023risk ; su2024optimality . A risk-averse policy prefers actions with more certainty even if it means a lower expected return shen2014risk ; mazouchi2023risk . The goal of Risk-averse RL (RARL) is to compute risk-averse policies. To this end, RARL specifies the objective using monetary risk measures, such as value-at-risk (VaR), conditional-value-at-risk (CVaR), entropic risk measure (ERM), or entropic value at risk (EVaR). These risk measures penalize the variability of returns interpretably and yield policies with stronger guarantees on the probability of catastrophic losses Follmer2016stochastic .
@@
This paper derives risk-averse TRC model-free Q-learning algorithms for ERM-TRC and EVaR-TRC objectives. There are three main challenges to deriving these risk-averse TRC Q-learning algorithms. First, generalizing the standard Q-learning algorithm to risk-averse objectives requires computing the full return distribution rather than just its expectation. Second, the risk-averse TRC Bellman operator may not be a contraction. The absence of contraction property precludes the direct application of model-free methods in an undiscounted MDP. Third, instead of the contraction property, we need to rely on other properties of the Bellman operator and an additional bounded condition to prove the convergence of the risk-averse TRC Q-learning algorithms.
As our main contribution, we derive a rigorous proof of the ERM-TRC Q-learning algorithm and the EVaR-TRC Q-learning algorithm converging to the optimal risk-averse value functions. We also show that the proposed Q-learning algorithms compute the optimal stationary policies for the ERM and EVaR objectives, and the optimal state-action value function $q$ can be computed by the stochastic gradient descent along the derivative of the exponential loss function.
The rest of the paper is organized as follows. We define our research setting and preliminary concepts in Section 2. In Section 3, we leverage the elicitability of ERM, define a new ERM Bellman operator, and propose Q-learning algorithms for the ERM and EVaR objectives. In Section 4, we give a rigorous convergence proof of the proposed ERM-TRC Q-learning and EVaR-TRC Q-learning algorithms. Finally, the numerical results presented in Section 5 illustrate the effectiveness of our algorithms.
@@ 2 Preliminaries for Risk-aversion in Markov Decision Processes
We denote by $\mathbb{R}$ and $\mathbb{N}$ the sets of real and natural (including 00) numbers, and $\bar{\mathbb{R}}:=\mathbb{R}\cup\left\{-\infty,\infty\right\}$ denotes the extended real line. We use $\mathbb{R}_{+}$ and $\mathbb{R}_{++}$ to denote non-negative and positive real numbers, respectively. We use a tilde to indicate a random variable, such as $\tilde{x}\colon\Omega\to\mathbb{R}$ for the sample space $\Omega$. The set of all real-valued random variables is denoted as $\mathbb{X}:=\mathbb{R}^{\Omega}$. Sets are denoted with calligraphic letters.
@@
$\operatorname{ERM}_{\beta}\left[\tilde{x}\right]\;:=\;-\beta^{-1}\cdot\log\mathbb{E}\operatorname{exp}\left(-\beta\cdot\tilde{x}\right),$ (1)
and can be extended to $\beta\in[0,\infty]$ as $\operatorname{ERM}_{0}[\tilde{x}]=\lim_{\beta\to 0^{+}}\operatorname{ERM}_{\beta}\left[\tilde{x}\right]=\mathbb{E}[\tilde{x}]$ and $\operatorname{ERM}_{\infty}[\tilde{x}]=\lim_{\beta\to\infty}\operatorname{ERM}_{\beta}\left[\tilde{x}\right]=\operatorname{ess}\inf[\tilde{x}]$. ERM is popular because of its simplicity and its favorable properties in multi-stage optimization formulations su2025risk ; hau2023entropic ; Hau2023a ; hau2024q . In particular, dynamic decision-making with ERM allows for the existence of dynamic programming equations and Markov or stationary optimal policies. In addition, in this work we leverage the fact that ERM is elicitable which means that it can be estimated by solving a linear regression problem bellini2015elicitable ; embrechts2021bayes .
The second risk measure we consider is entropic value at risk (EVaR), which is defined for a given risk level $\alpha\in(0,1)$ and $\tilde{x}\in\mathbb{X}$ as
$\operatorname{EVaR}_{\alpha}\left[\tilde{x}\right]$ $\;:=\;\sup_{\beta>0}-\beta^{-1}\log\left(\alpha^{-1}\mathbb{E}\operatorname{exp}\left(-\beta\tilde{x}\right)\right)\;=\;\sup_{\beta>0}\operatorname{ERM}_{\beta}\left[\tilde{x}\right]+\beta^{-1}\log\alpha,$ (2)
@@
Recent work has shown that an optimal policy in (3) exists, is stationary, and has bounded return for a sufficiently small $\beta$ as long as the following assumptions hold su2025risk . The sink state $e$ satisfies that $p(e,a,e)=1$ and $r(e,a,e)=0$ for each $a\in\mathcal{A}$, and $\mu_{e}=0$ su2025risk . It is crucial to assume that the MDP is transient for any $\pi\in\Pi$:
@@
With a risk-neutral objective, transient MDPs guarantee that the return is bounded for each policy. However, that is no longer the case with the ERM objective. In particular, for some $\beta$ it is possible that the return is unbounded. The return is bounded for a sufficiently small $\beta$ and there exists an optimal stationary policy su2025risk . In contrast, the return of EVaR is always bounded, and an optimal stationary policy always exists regardless of the risk level $\alpha$ su2025risk .
@@
$\tilde{q}_{i+1}(\tilde{s}_{i},\tilde{a}_{i})=\tilde{q}_{i}(\tilde{s}_{i},\tilde{a}_{i})-\tilde{\eta}_{i}\tilde{z}_{i},\quad\tilde{z}_{i}=r(\tilde{s}_{i},\tilde{a}_{i},\tilde{s}^{\prime}_{i})+\max_{a^{\prime}\in\mathcal{A}}\tilde{q}_{i}(\tilde{s}_{i},a^{\prime})-\tilde{q}_{i}(\tilde{s}_{i},\tilde{a}_{i}).$ (5)
Here, $\tilde{z}_{i}$ is also known as the TD residual. The algorithm assumes a stream of samples $(\tilde{s}_{i},\tilde{a}_{i},\tilde{s}^{\prime}_{i})_{i=1,\dots}$ sampled from the transition probabilities and appropriately chosen step sizes $\tilde{\eta}_{i}$ to converge to the optimal state-action value function. Note that $\tilde{q}_{i}$ is random because it is a function of a random variable. It is well-known that the standard Q-learning algorithm can be seen as a stochastic gradient descent on the quadratic loss function hau2024q ; Asadi2024 ; shen2014risk . We leverage this property to build our algorithms.
@@ 3 Q-learning Algorithms: ERM and EVaR
In this section, we derive new Q-learning algorithms for the ERM and EVaR objectives. First, we propose the Q-learning algorithm for ERM in Section 3.1, which requires us to introduce a new ERM Bellman operator based on the elicitability of the ERM risk measure. Then, we use this algorithm in Section 3.2 to propose an EVaR Q-learning algorithm. The proofs for this section are deferred to Appendix A.
@@ 3.1 ERM Q-learning Algorithm
The algorithm we propose in this section computes the state-action function for multiple values of the risk level $\beta\in\mathcal{B}$ for some given non-empty finite set $\mathcal{B}\subseteq\mathbb{R}_{++}$ of positive reals. The set $\mathcal{B}$ may be a singleton or may include multiple values, which is necessary to optimize EVaR in the next section.
@@
The superscript $s,a\right>$ in (6) indicates the initial state and action are $\tilde{s}_{0}=s,\tilde{a}_{0}=a$, and the policy $\pi$ determines the actions henceforth. We use $\mathcal{Q}:=\mathbb{R}^{\mathcal{S}\times\mathcal{A}}$ to denote the set of possible state-action value functions.
@@
The following result, which follows from (su2025risk, , theorem 3.3), shows that the ERM state-action value function can be computed as the fixed point of the Bellman operator.
@@
We now turn to an alternative definition of the Bellman operator that can be used to estimate the value function directly from samples. To develop our ERM Q-learning algorithm, we need to define a Bellman operator that is amenable to compute its fixed point using stochastic gradient descent. For this purpose, we use the elicitability property of risk measures bellini2015elicitable . ERM is known to be elicitable using the following loss function $\ell_{\beta}\colon\mathbb{R}\to\mathbb{R}$:
@@
The functions $\ell^{\prime}$ and $\ell^{\prime\prime}$ are the first and second derivatives which are be important in constructing and analyzing the Q-learning algorithm.
@@
Following Proposition 3.2, we get the following equivalence of the Bellman operator, which implies that their fixed points coincide.
@@
We can introduce the ERM Q-learning algorithm in Algorithm 1 using the results above. This algorithm adapts the standard Q-learning approach to the risk-averse setting. Intuitively, it works as follows. Each iteration $i$ processes a single transition sample to update the optimal state-action value function estimate $\tilde{q}_{i}$. The value function estimates are random because the samples are random. The value function estimates are updated following a stochastic gradient descent along the derivative of the loss function $\ell_{\beta}$ defined in (9). Each sample is used to simultaneously update the state-action value function for multiple values of $\beta\in\mathcal{B}$.
Input: Risk levels $\mathcal{B}\subseteq\mathbb{R}$, samples: $(\tilde{s}_{i},\tilde{a}_{i},\tilde{s}_{i}^{\prime})$, step sizes $\tilde{\eta}_{i},i\in\mathbb{N}$, bounds $z_{\min},z_{\max}$
@@ 1 $\tilde{q}_{0}(s,a,\beta)\leftarrow 0,\quad\forall s\in\mathcal{S},a\in\mathcal{A}$ ;
5             if $\neg(z_{\min}\leq\tilde{z}_{i}(\beta)\leq z_{\max})$ then return $\tilde{q}=-\infty$;
@@
There are three main differences between Algorithm 1 and the standard Q-learning algorithm described in Section 2. First, standard Q-learning aims to maximize the expectation objective, and the proposed algorithm maximizes the ERM-TRC objective. Second, the state-action value function update follows a sequence of stochastic gradient steps, but it replaces the quadratic loss function with the exponential loss function $\ell_{\beta}$ in (9). Third, $q$ values in the proposed algorithm need an additional bounded condition that the TD residual $\tilde{z}_{i}(\beta)$ is in $[z_{\min},z_{\max}]$.
@@
The $q$ value can be unbounded for two main reasons. First, as the value of $\beta$ increases, the ERM value function does not increase (hau2023entropic, , Lemma A.7) and can reach $\infty$- ∞ (su2025risk, , Theorem 3.3). Second, the agent has some probability of repeatedly receiving a reward for the same action, leading to an uncontrolled increase in the $q$ value for that action.
Note that if $\tilde{z}_{i}(\beta)$ is outside the $[z_{\min},z_{\max}]$ range, it indicates that the value of the risk level $\beta$ is so large that $\tilde{q}$ is unbounded. This aligns with the conclusion that the value function may be unbounded for a large risk factor su2025risk ; patek2001terminating . Detecting unbounded values of $q$ caused by random chance is useful but is challenging and beyond the scope of this work.
@@
We now discuss how to estimate $z_{\min}$ and $z_{\max}$ in Algorithm 1. Assume that the sequences $\{\tilde{\eta}_{i}\}_{i=0}^{\infty}$ and $\{(\tilde{s}_{i},\tilde{a}_{i},\tilde{s}^{\prime}_{i})\}_{i=0}^{\infty}$ used in Algorithm 1, $\beta>0$, we have
@@
Where $\|r\|_{\infty}=\max_{s,s^{\prime}\in\mathcal{S},a\in\mathcal{A}}|r(s,a,s^{\prime})|$. The constant $c$ and $d$ can be estimated by Algorithm 3 in the appendix. For more details on the derivation, we refer the interested reader to Section A.3.

### v2

@@ 1 Introduction
Risk-averse reinforcement learning (RL) is an essential framework for practical and high-stakes applications, such as self-driving, robotic surgery, healthcare, and finance greenberg2022efficient ; fei2021exponential ; urpi2021risk ; hau2024q ; suevar ; zhang2021mean ; hau2023entropic ; Kastner2023 ; Marthe2023a ; lam2022risk ; li2022quantile ; bauerle2022markov ; pan2019risk ; mazouchi2023risk ; su2024optimality . A risk-averse policy prefers actions with more certainty even if it means a lower expected return shen2014risk ; mazouchi2023risk . The goal of risk-averse RL is to compute risk-averse policies. To this end, risk-averse RL specifies the objective using monetary risk measures, such as value-at-risk (VaR), conditional value-at-risk (CVaR), entropic risk measure (ERM), or entropic value-at-risk (EVaR). These risk measures penalize the variability of returns interpretably and yield policies with stronger guarantees on the probability of catastrophic losses Follmer2016stochastic .
@@
This paper derives risk-averse TRC model-free Q-learning algorithms for ERM-TRC and EVaR-TRC objectives. There are three main challenges to deriving these risk-averse TRC Q-learning algorithms. First, generalizing the standard Q-learning algorithm to risk-averse objectives requires computing the full return distribution rather than just its expectation. Second, the risk-averse TRC Bellman operator may not be a contraction. The absence of a contraction property precludes the direct application of model-free methods in an undiscounted MDP. Third, instead of the contraction property, we need to rely on other properties of the Bellman operator and an additional bounded condition to prove the convergence of the risk-averse TRC Q-learning algorithms.
As our main contribution, we derive a rigorous proof of the ERM-TRC Q-learning algorithm and the EVaR-TRC Q-learning algorithm converging to the optimal risk-averse value functions. We also show that the proposed Q-learning algorithms compute the optimal stationary policies for the ERM-TRC and EVaR-TRC objectives, and the optimal state-action value function can be computed by stochastic gradient descent along the derivative of the exponential loss function.
The rest of the paper is organized as follows. We define our research setting and preliminary concepts in Section˜2. In Section˜3, we leverage the elicitability of ERM, define a new ERM Bellman operator, and propose Q-learning algorithms for the ERM and EVaR objectives. In Section˜4, we give a rigorous convergence proof of the proposed ERM-TRC Q-learning and EVaR-TRC Q-learning algorithms. Finally, the numerical results presented in Section˜5 illustrate the effectiveness of our algorithms.
@@ 2 Preliminaries for Risk-aversion in Markov Decision Processes
We denote by $\mathbb{R}$ and $\mathbb{N}$ the sets of real and natural (including 0) numbers, and $\bar{\mathbb{R}}:=\mathbb{R}\cup\left\{-\infty,\infty\right\$} denotes the extended real line. We use $\mathbb{R}_{+}$ and $\mathbb{R}_{++}$ to denote non-negative and positive real numbers, respectively. We use a tilde to indicate a random variable, such as $\tilde{x}\colon\Omega\to\mathbb{R}$ for the sample space $\Omega$. The set of all real-valued random variables is denoted as $\mathbb{X}:=\mathbb{R}^{\Omega}$. Sets are denoted with calligraphic letters.
@@
[…]
and can be extended to $\beta\in[0,\infty]$ as $\operatorname{ERM}_{0}[\tilde{x}]=\lim_{\beta\to 0^{+}}\operatorname{ERM}_{\beta}\left[\tilde{x}\right]=\mathbb{E}[\tilde{x}$] and $\operatorname{ERM}_{\infty}[\tilde{x}]=\lim_{\beta\to\infty}\operatorname{ERM}_{\beta}\left[\tilde{x}\right]=\operatorname{ess}\inf[\tilde{x}$]. ERM is popular because of its simplicity and its favorable properties in multi-stage optimization formulations su2025risk ; hau2023entropic ; Hau2023a ; hau2024q . In particular, dynamic decision-making with ERM allows for the existence of dynamic programming equations and Markov or stationary optimal policies. In addition, in this work, we leverage the fact that ERM is elicitable, which means that it can be estimated by solving a linear regression problem bellini2015elicitable ; embrechts2021bayes .
The second risk measure we consider is entropic value-at-risk (EVaR), which is defined for a given risk level $\alpha\in(0,1)$ and $\tilde{x}\in\mathbb{X}$ as
$\operatorname{EVaR}_{\alpha}\left[\tilde{x}\right$] $\;=\;\sup_{\beta>0}-\beta^{-1}\log\left(\alpha^{-1}\mathbb{E}\operatorname{exp}\left(-\beta\tilde{x}\right)\right)\;=\;\sup_{\beta>0}\operatorname{ERM}_{\beta}\left[\tilde{x}\right]+\beta^{-1}\log\alpha,$ (2)
@@
Recent work has shown that an optimal policy in (3) exists, is stationary, and has bounded return for a sufficiently small $\beta$ as long as the following assumptions hold su2025risk . The sink state $e$ satisfies that $p(e,a,e)=1$ and $r(e,a,e)=0$ for each $a\in\mathcal{A}$, and $\mu_{e}=0$. It is crucial to assume that the MDP is transient for any $\pi\in\Pi$:
@@
With a risk-neutral objective, transient MDPs guarantee that the return is bounded for each policy. However, that is no longer the case with the ERM objective. In particular, for some $\beta$ it is possible that the return is unbounded. The return is bounded for a sufficiently small $\beta$ and there exists an optimal stationary policy su2025risk . In contrast, the return of EVaR is always bounded, and an optimal stationary policy always exists regardless of the risk level $\alpha$ su2025efficient .
@@
$\tilde{q}_{i+1}(\tilde{s}_{i},\tilde{a}_{i})=\tilde{q}_{i}(\tilde{s}_{i},\tilde{a}_{i})-\tilde{\eta}_{i}\tilde{z}_{i},\quad\tilde{z}_{i}=r(\tilde{s}_{i},\tilde{a}_{i},\tilde{s}^{\prime}_{i})+\max_{a^{\prime}\in\mathcal{A}}\tilde{q}_{i}(\tilde{s}^{\prime}_{i},a^{\prime})-\tilde{q}_{i}(\tilde{s}_{i},\tilde{a}_{i}).$ (5)
Here, $\tilde{z}_{i}$ is also known as the TD residual. The algorithm assumes a stream of samples $(\tilde{s}_{i},\tilde{a}_{i},\tilde{s}^{\prime}_{i})_{i=1,\dots}$ sampled from the transition probabilities and appropriately chosen step sizes $\tilde{\eta}_{i}$ to converge to the optimal state-action value function. Note that $\tilde{q}_{i}$ is random because it is a function of a random variable. We restate lemma C.13 from hau2024q as the following lemma because it shows the connection between Q-learning and gradient descent.
Lemma 2.1 (Lemma C.13 in hau2024q ).
Suppose that $f\colon\mathbb{R}\rightarrow\mathbb{R}$ is a differentiable $\mu$-strongly convex function with an $L$-Lipschitz continuous gradient. Consider $x_{i}\in\mathbb{R}$ and a gradient update for any step size $\xi\in(0,1/L]:$
$x_{i+1}:=x_{i}-\xi\cdot f^{\prime}(x_{i}).$
Then $\exists l\in[1/L,1/\mu]$ such that $\xi/l\in(0,1]$ and
$x_{i+1}$ $=(1-\xi/l)\cdot x_{i}+\xi/l\cdot x^{\star},$
where $x^{\star}=\operatorname*{arg\,min}_{x\in\mathbb{R}}f(x)$ is unique from the strong convexity of $f$.
The standard Q-learning algorithm can be seen as a stochastic gradient descent on the quadratic loss function $f$ hau2024q ; Asadi2024 ; shen2014risk . We leverage this property to build our algorithms.
@@ 3 Q-learning Algorithms: ERM and EVaR
In this section, we derive new Q-learning algorithms for the ERM and EVaR objectives. First, we propose the Q-learning algorithm for ERM in Section˜3.1, which requires us to introduce a new ERM Bellman operator based on the elicitability of the ERM risk measure. Then, we use this algorithm in Section˜3.2 to propose an EVaR Q-learning algorithm. The proofs for this section are deferred to Appendix˜A.
@@ 3.1 ERM Q-learning Algorithm
The algorithm we propose in this section computes the state-action value function for multiple values of the risk level $\beta\in\mathcal{B}$ for some given non-empty finite set $\mathcal{B}\subseteq\mathbb{R}_{++}$. The set $\mathcal{B}$ may be a singleton or may include multiple values, which is necessary to optimize EVaR in the next section.
@@
The superscript $\left<s,a\right$> in (6) indicates the initial state and action are $\tilde{s}_{0}=s,\tilde{a}_{0}=a$, and the policy $\pi$ determines the actions henceforth. We use $\mathcal{Q}:=\mathbb{R}^{\mathcal{S}\times\mathcal{A}}$ to denote the set of possible state-action value functions.
@@
The following result, which follows from Theorem˜A.1, shows that the ERM state-action value function can be computed as the fixed point of the Bellman operator.
@@
We now turn to an alternative definition of the Bellman operator that can be used to estimate the value function directly from samples. To develop our ERM Q-learning algorithm, we need to define a Bellman operator that is amenable to computing its fixed point using stochastic gradient descent. For this purpose, we use the elicitability property of risk measures bellini2015elicitable . ERM is known to be elicitable using the following loss function $\ell_{\beta}\colon\mathbb{R}\to\mathbb{R}:$
@@
The functions $\ell^{\prime}$ and $\ell^{\prime\prime}$ are the first and second derivatives, which are important in constructing and analyzing the Q-learning algorithm.
@@
Following Proposition˜3.2, we get the following equivalence of the Bellman operator, which implies that their fixed points coincide.
@@
We can introduce the ERM Q-learning algorithm in Algorithm˜1 using the results above. This algorithm adapts the standard Q-learning approach to the risk-averse setting. Intuitively, it works as follows. Each iteration $i$ processes a single transition sample to update the optimal state-action value function estimate $\tilde{q}_{i}$. The value function estimates are random because the samples are random. The value function estimates are updated following a stochastic gradient descent along the derivative of the loss function $\ell_{\beta}$ defined in (9). Each sample is used to simultaneously update the state-action value function for multiple values of $\beta\in\mathcal{B}$.
Input: Risk levels $\mathcal{B}\subseteq\mathbb{R}_{++}$, samples: $(\tilde{s}_{i},\tilde{a}_{i},\tilde{s}_{i}^{\prime})$, step sizes $\tilde{\eta}_{i},i\in\mathbb{N}$, bounds $z_{\min},z_{\max}$
@@ 1 $\tilde{q}_{0}(s,a,\beta)\leftarrow 0,\quad\forall s\in\mathcal{S},a\in\mathcal{A}$ ;
5    if $\neg(z_{\min}\leq\tilde{z}_{i}(\beta)\leq z_{\max})$ then return $\tilde{q}_{i+1}(s,a,\beta)=$-$\infty$;
@@
There are three main differences between Algorithm˜1 and the standard Q-learning algorithm described in Section˜2. First, standard Q-learning aims to maximize the expectation objective, and the proposed algorithm maximizes the ERM-TRC objective. Second, the state-action value function update follows a sequence of stochastic gradient steps, but it replaces the quadratic loss function with the exponential loss function $\ell_{\beta}$ in (9). Third, $q$ values in the proposed algorithm need an additional bounded condition that the TD residual $\tilde{z}_{i}(\beta)$ is in $z_{\min},z_{\max}]$.
@@
The $q$ value can be unbounded for two main reasons. First, as the value of $\beta$ increases, the ERM value function does not increase, as shown in Lemma˜A.2, and can reach -$\infty$, as shown in Theorem˜A.1. Second, the agent has some probability of repeatedly receiving a reward for the same action, leading to an uncontrolled increase in the $q$ value for that action.
Note that if $\tilde{z}_{i}(\beta)$ is outside the $z_{\min},z_{\max}]$ range, it indicates that the value of the risk level $\beta$ is so large that $\tilde{q}$ is unbounded. This aligns with the conclusion that the value function may be unbounded for a large risk factor su2025risk ; patek2001terminating . Detecting unbounded values of $q$ caused by random chance is useful but is challenging and beyond the scope of this work.
@@
We now discuss how to estimate $z_{\min}$ and $z_{\max}$ in Algorithm˜1. Assume that the sequences $\{\tilde{\eta}_{i}\}_{i=0}^{\infty}$ and $\{(\tilde{s}_{i},\tilde{a}_{i},\tilde{s}^{\prime}_{i})\}_{i=0}^{\infty}$ used in Algorithm˜1, $\beta>0$, we have
@@
Where $\|r\|_{\infty}=\max_{s,s^{\prime}\in\mathcal{S},a\in\mathcal{A}}|r(s,a,s^{\prime})|$. The constants $c$ and $d$ can be estimated by Algorithm˜3 in the appendix. For more details on the derivation, we refer the interested reader to Section˜A.4.
Now we explain how to leverage the elicitability and monotonicity of ERM to design Algorithm˜1 as a stochastic gradient descent. Let $b=(s,a,\beta),s\in\mathcal{S},a\in\mathcal{A}$, $\beta\in\mathcal{B}$, and $\xi>0$, and define operators $G$ and $H$ as
$(Gq)(b)$ $=\frac{\partial}{\partial y}\mathbb{E}^{a,s}\Bigl[\ell_{\beta}\Bigl(r(s,a,\tilde{s}_{1})+\max_{a^{\prime}\in\mathcal{A}}q(\tilde{s}_{1},a^{\prime},\beta)-y\Bigr)\Bigr]\mid_{y=q(s,a,\beta)},$ (12)
$(G_{s^{\prime}}q)(b)$ $=\frac{\partial}{\partial y}\mathbb{E}\Bigl[\ell_{\beta}\Bigl(r(s,a,s^{\prime})+\max_{a^{\prime}\in\mathcal{A}}q(s^{\prime},a^{\prime},\beta)-y\Bigr)\Bigr]\mid_{y=q(s,a,\beta)}$ (13)
The operation $(Gq)(b)$ can be interpreted as the average gradient step where $\tilde{s}_{1}$ is a random variable representing the next state. The operation $(G_{s^{\prime}}q)(b)$ can be interpreted as the individual gradient step for a specific next state $s^{\prime}$.
$(Hq)(b):=q(b)-\xi\cdot(Gq)(b),\qquad(H_{s^{\prime}}q)(b):=q(b)-\xi\cdot(G_{s^{\prime}}q)(b).$ (14)
The operation $(Hq)(b)$ can be interpreted as the q-update for a random variable $\tilde{s}_{1}$ representing the next state. $(H_{s^{\prime}}q)(b)$ can be interpreted as the q-update a specific next state $s^{\prime}$.
Lemma 3.4.
$x(b)\leq y(b)$ $\Rightarrow(Hx)(b)\leq(Hy)(b)$ (a)
$(Hq^{\star})(b)$ $=q^{\star}(b)$ (b)
The proof has two main steps. First, rewrite $(Hq)(b)$ in terms of the ERM Bellman operator in (11). Second, prove the monotonicity property of $(Hq)(b)$, which is necessary for the convergence analysis of Algorithm˜1. Fix $\beta>0$ and some $b=(s,a,\beta),s\in\mathcal{S},a\in\mathcal{A}$, fix $q$ and define
$f(y)=\mathbb{E}^{s,a}\bigl[\ell_{\beta}(r(s,a,\tilde{s}_{1})+\max_{a^{\prime}\in\mathcal{A}}q(\tilde{s}_{1},a^{\prime},\beta)-y)\bigr].$ (15)
The function $f$ is $\ell$-strongly convex with an $L$-Lipschitz-continuous gradient from Lemma˜4.4. Let $y^{\star}=\arg\min_{y\in\mathbb{R}}f(y)=(\hat{B}q)(b)$ and $\exists l\in[1/L,1/\ell$] such that
$(Hq)(b)$ $=(q-\xi Gq)(b)\stackrel{{\scriptstyle\text{(a)}}}{{=}}q(b)-\xi f^{\prime}(q(b))\stackrel{{\scriptstyle\text{(b)}}}{{=}}(1-\xi/l)q(b)+\xi/l\cdot y^{\star}$ (16)
$\stackrel{{\scriptstyle\text{(c)}}}{{=}}(1-\xi/l)q(b)+\xi/l\cdot(\hat{B}q)(b).$
Step (a) follows by algebraic manipulation from the definitions in (12) and (15). Step (b) follows from Lemma˜2.1. Step (c) follows from the fact that $\hat{B}$ is the ERM Bellman operator defined in (11) and $y^{\star}$ is the unique solution to (10).